# Bi-level regularization via iterative mesh refinement for aeroacoustics

### v2

@@ Abstract
In this work, we illustrate the connection between adaptive mesh refinement for finite element discretized PDEs and the recently developed bi-level regularization algorithm. By adaptive mesh refinement according to data noise, regularization effect and convergence are immediate consequences. We moreover demonstrate its numerical advantages to the classical Landweber algorithm in term of time and reconstruction quality for the example of the Helmholtz equation in an aeroacoustic setting.
@@
inverse source problem, bi-level approach, Landweber method, Helmholtz equation, finite element, mesh refinement.
@@ 0.1 Inverse aeroacoustic source with Helmholtz equation
$\phi\in L^{2}(\Omega_{0})\qquad\Omega_{0}\subset\text{int }\Omega$ (1)
@@
$\begin{split}M:&\,H^{1}(\Omega\setminus\bigcup_{i=1}^{3}S_{i},\mathbb{C})\to L^{2}(\Omega_{1},\mathbb{C})\qquad u\mapsto u|_{\Omega_{1}}\end{split}$ (2)
Acoustic wave traveling in the room at a constant wave number $k\in\mathbb{R}$ is governed by the homogeneous Helmholtz equation
@@
where $\phi$ is implicitly extended by zero outside of $\Omega_{0}$. The sound-hard scatterers are modeled by the Neumann condition on the interior Lipschitz boundaries $\partial S_{i}$, and on the exterior boundary $\partial\Omega$, we use an approximate Sommerfeld radiation condition; $n$ are the outward normal vectors.
@@
$\text{Find }\phi:\quad F(\phi)=y^{\delta}.$
@@
We shall use the bi-level regularization scheme (nguyen24, , Algorithm 1.2) for the reconstruction process. Here, the author proposes a novel bi-level inversion framework, in which the upper level-iteration for the unknown parameter – here, the source term – embeds a lower-iteration solving nonlinear PDEs inexactly. A key contribution of nguyen24 is the derivation of stopping rules for upper- and lower-level that can ensure convergence. In addition, embedding any PDE solver (FD, FEM, multigrid etc.) into the lower-level, the error analysis in nguyen24 can inform an adaptive discretization scheme via multi-grid size, -mesh size or -resolution Trottenberg that are computationally effective by not fixing a single fine scale in the inversion process.
As the model (3) is rather classical, we shall not use the lower-level to approximate $u$, instead implementing a FEM solver with an adaptive mesh refinement strategy suggested in (nguyen24, , Theorem 2) to speed up the reconstruction, while, crucially, maintaining the regularizing property. This forms the contribution of this article.
@@ 0.2 Inversion algorithm with mesh iterative refinement
where $\overline{(\cdot)}$ denotes the complex conjugate. The proof is complete.
@@ 0.3 Numerical experiments
$\phi(x,y):=\sqrt{\min\{\tfrac{1}{4}-x^{2}-y^{2},0\}}\cos\left(2\pi\sqrt{x^{2}+y^{2}}\right).$
@@
For the approximate PDE solver, we employ the NGSolve Schoberl finite element method (FEM) package to create a discretization of our domain $\Omega:=[-1,1]^{2}\subset\mathbb{R}^{2}$ and subsequently also the forward operator $F$, where we have chosen $\Omega_{0}:=[-1/2,1/2]^{2}\subset\Omega$ and $\Omega_{1}:=[-1,1]^{2}\setminus[-11/20,11/20]^{2}\subset\mathbb{R}^{2}$ (see Figures; sound-hard scatterers $S_{i}$, $i=1,2,3$ are as modeled). We approximate the true state $u:=F\phi$ by applying a finely discretized FEM solver ($h^{\dagger}\approx 0.046$), and conduct two numerical comparisons for the recovery of $\phi$ from noisy observations $y^{\delta}$, where $y^{\delta}-u$ consists of additive Gaussian white noise $\delta$ of relative size $1\%$ resp. $10\%$ relative to $u$. In both experiments, two methods of recovering $\phi$ are compared.
Figure 2: Left: Output of bi-level. Right: Output of direct Landweber. Bottom: True source $\phi$. $1\%$ noise, mesh not depicted.
The first method is the bi-level Landweber iteration with mesh refinement as proposed above, beginning with a coarse mesh ($h_{0}\approx 0.531$), $q:=2^{1/60}$ and $C:=1.4/\delta$. The second method is “direct” Landweber regularization, carried out on a single, distinct FEM-discretized mesh that is finer than any used by the bi-level and coarser than the FEM used to generate the approximate state $u$ ($h_{\text{direct}}\approx 0.064$). In all cases, $\mu$ was estimated as $\mu\approx 0.075$, and all iteration was stopped at the first (upper-)index $j$ satisfying the discrepancy principle $\|y^{\delta}-F\phi^{j}\|_{2}\leq\tau\delta$ with $\tau:=1.3$.
Figure 3: Iterative reconstruction on each mesh in the bi-level algorithm, prior to refinement. Noise level $1\%$. $h\approx 0.531$, $0.265$.
We make some remarks regarding the reconstruction. Figure 4 displays the reconstruction errors and residuals for the two methods, plotted by total time taken. This includes time required for mesh refinement (bi-level only), for the Landweber step and for computation of the residual.
@@
Conclusions
Figure 5: Left: Output of bi-level. Right: Output of direct Landweber. Bottom: True source $\phi$. $10\%$ noise, mesh not depicted.
@@ Conclusions
The recent developments in bi-level algorithms nguyen24 ; nguyen-seqbi opens up exciting usage within the field of optimal experimental design (OED). Aarset newly proposed a novel, high-speed algorithm for linear OED. As remarked in Ucinski , detailed treatment of the non-linear case would require additional developments within bi-level algorithms, which the authors believe may now be available.

### v3

@@ Abstract
In this work, we illustrate the connection between adaptive mesh refinement for finite element discretized PDEs and the recently developed bi-level regularization algorithm, due to Nguyen (Inverse Problems 40(4) 2024). By adaptive mesh refinement according to data noise, regularization effect and convergence are immediate consequences. We moreover demonstrate its numerical advantages to the classical Landweber algorithm in terms of time and reconstruction quality for the example of the Helmholtz equation in an aeroacoustic setting.
@@
inverse source problem, bi-level approach, Landweber method, Helmholtz equation, finite elements, mesh refinement.
@@ 0.1 Inverse aeroacoustic source with Helmholtz equation
$\phi\in L^{2}(\Omega_{0}),\quad\Omega_{0}\subset\text{int }\Omega$ (1)
@@
$\begin{split}M:&\,H^{1}(\Omega\setminus\bigcup_{i=1}^{3}S_{i},\mathbb{C})\to L^{2}(\Omega_{1},\mathbb{C}),\quad u\mapsto u|_{\Omega_{1}}.\end{split}$ (2)
The travel of acoustic waves through the room is for each constant wave number $k\in\mathbb{R}$ governed by the homogeneous Helmholtz equation
@@
where $\phi$ is implicitly extended by zero outside of $\Omega_{0}$. The sound-hard scatterers are modeled by the Neumann condition on the interior Lipschitz boundaries $\partial S_{i}$, and on the exterior boundary $\partial\Omega$ we use an approximate Sommerfeld radiation condition; $n$ are the outward normal vectors.
@@
Find ϕ such $\text{Find $\phi$ such that}\quad F(\phi)=y^{\delta}.$
@@
We shall use the bi-level regularization scheme (nguyen24, , Algorithm 1.2) for the reconstruction process. In the cited work, the author proposes a novel bi-level inversion framework. The upper level-iteration, which iteratively approximates the unknown parameter – for this article, the source term – embeds a lower level-iteration, which solves nonlinear PDEs inexactly. A key contribution of nguyen24 is the derivation of stopping rules for upper- and lower-level that can ensure convergence. In addition, embedding any PDE solver (FD, FEM, multigrid etc.) into the lower-level, the error analysis in nguyen24 can inform an adaptive discretization scheme via multi-grid size, -mesh size or -resolution Trottenberg that are computationally effective by not fixing a single fine scale in the inversion process.
As the model (3) is rather classical, we shall not use the lower-level to approximate $u$. Instead, Section 0.2 details how to implement a FEM solver with an adaptive mesh refinement strategy suggested in (nguyen24, , Theorem 2) to speed up the reconstruction, while, crucially, maintaining the regularizing property. This, along with the numerical verification provided in Section 0.3, forms the contribution of this article.
@@ 0.2 Inversion algorithm with mesh iterative refinement
where $\overline{(\cdot)}$ denotes the complex conjugate, $u$ is any solution of (3) on $\Omega\setminus\bigcup_{i=1}^{3}S_{i}$, and recalling that $v$ is extended by zero outside of $\Omega_{1}$. The proof is complete.
@@ 0.3 Numerical experiments
$\phi(x_{1},x_{2}):=\sqrt{\min\{\tfrac{1}{4}-x_{1}^{2}-x_{2}^{2},0\}}\cos\left(2\pi\sqrt{x_{1}^{2}+x_{2}^{2}}\right).$
For the approximate PDE solver, we employ the NGSolve Schoberl finite element method (FEM) package to discretize our domain $\Omega:=[-1,1]^{2}\subset\mathbb{R}^{2}$ and subsequently also the forward operator $F$, where $\Omega_{0}:=[-1/2,1/2]^{2}\subset\Omega$ and $\Omega_{1}:=[-1,1]^{2}\setminus[-11/20,11/20]^{2}\subset\mathbb{R}^{2}$ (see Figures; sound-hard scatterers $S_{i}$, $i=1,2,3$ are as modeled). We approximate the true state $u:=F\phi$ by applying a finely discretized FEM solver ($h^{\dagger}\approx 0.046$), and conduct two numerical experiments, comparing the recovery of the source term $\phi$ from noisy observations $y^{\delta}$ of the true state $u$, where $y^{\delta}-u$ consists of additive Gaussian white noise of size $1\%$ resp. $10\%$ relative to $u$. In both experiments, two methods of recovering $\phi$ are compared.
The first method is bi-level Landweber iteration with mesh refinement as proposed above, beginning with a coarse mesh ($h_{0}\approx 0.531$), $q:=2^{1/60}$ and $C:=1.4/\delta$. The second method is “direct” Landweber regularization, carried out on a single, distinct FEM-discretized mesh that is finer than any used by the bi-level and coarser than the FEM used to generate the approximate state $u$ ($h_{\text{direct}}\approx 0.064$). In all cases, $\mu$ was estimated as $\mu\approx 0.075$, and all iteration was stopped at the first (upper-)index $j$ satisfying the discrepancy principle $\|y^{\delta}-F\phi^{j}\|_{2}\leq\tau\delta$ with $\tau:=1.3$.
@@
Figure 2: Iterative reconstruction on each mesh in the bi-level algorithm, prior to refinement. Noise level $10\%$. $h\approx 0.531$, $0.265$.
Figure 3: True source $\phi$. Mesh not depicted.
We make some remarks regarding the reconstruction. Figure 4 compares final reconstruction errors, while Figure 5 displays reconstruction errors and residuals for the two methods, plotted by total time taken, accounting for mesh refinement (bi-level only), the Landweber step and computation of the residual.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Error plots of reconstructions. Left to right: Bi-level with $1\%$ noise, direct Landweber with $1\%$ noise, bi-level with $10\%$ noise, direct Landweber with $10\%$ noise. Mesh not depicted.
@@
0.4 Conclusions
@@
The recent developments in bi-level algorithms nguyen24 ; nguyen-seqbi opens up exciting usage within the field of optimal experimental design (OED). The article Aarset newly proposed a novel method of OED for linear inverse problems. As remarked in Ucinski , detailed treatment of the non-linear case would require additional developments within bi-level algorithms, which the authors believe may now be available.